# 3INFN, Laboratori Nazionali del Gran Sasso, 67100 Assergi, Italy 4Max-Planck-Institut für Astronomie, Königstuhl 17, 69117, Heidelberg, Germany 5Universität Heidelberg, Zentrum für Astronomie (ZAH), Institut für Theoretische Astrophysik, Albert-Ueberle-Str. 2, 69120, Heidelberg, Germany

### v1

@@
[…]
@@
Very massive stars (VMSs, $M_{\star}$geq 100 M) play a crucial role in several astrophysical processes. At low metallicity, they might collapse directly into black holes, or end their lives as pair-instability supernovae. Recent observational results set an upper limit of $0.7\,{}\mathrm{yr}^{-1}\,{}\mathrm{Gpc}^{-3}$ on the rate density of pair-instability supernovae in the nearby Universe. However, most theoretical models predict rates exceeding this limit. Here, we compute new VMS tracks with the mesa code, and use them to analyze the evolution of the (pulsational) pair-instability supernova rate density across cosmic time. We show that stellar wind models accounting for the transition between optically thin and thick winds yield a pair-instability supernova rate $\mathcal{R}_{\mathrm{PISN}}\sim{}0.1$ Gpc-3 yr-1 at redshift $z\sim{}0$, about two orders of magnitude lower than our previous models. We find that the main contribution to the pair-instability supernova rate comes from stars with metallicity $Z\sim{}0.001-0.002$. Stars with higher metallicities cannot enter the pair-instability supernova regime, even if their zero-age main sequence mass is up to 500 M. The main reason is that VMSs enter the regime for optically thick winds during the main sequence at metallicity as low as $Z\sim{4}\times{}10^{-4}$. This enhances the mass loss rate, quenching the growth of the He core and thus preventing the onset of pair-instability in later evolutionary stages. This result highlights the critical role of mass loss in shaping the final fate of very massive stars and the rate of pair-instability supernovae.
@@ 1 Introduction
Very massive stars (VMSs), with an initial mass $>100\text{ M}_{\odot}$, represent a challenge to our understanding of stellar evolution due to their extreme physical properties. Several studies have explored their evolutionary paths (Belkus et al. 2007; Yungelson et al. 2008; Gräfener et al. 2011; Brott et al. 2011; Georgy et al. 2013; Yusof et al. 2013; Hirschi 2015; Limongi & Chieffi 2018) and their observational properties in the local Universe (Crowther et al. 2010; Bestenlehner et al. 2014).
As VMSs evolve, they lose mass via stellar winds, including nitrogen and other products of hot hydrogen burning. Hence, they might be one of the main sources of the N enhancement observed in some compact high-redshift galaxies (Vink 2023; Cameron et al. 2023; Charbonnel et al. 2023; Gieles et al. 2025). Several works show that, if they directly collapse onto a black hole, VMSs can form intermediate-mass black holes (IMBHs, with mass $>10^{2}$ M, Ebisuzaki et al. 2001; Heger & Woosley 2002; Portegies Zwart & McMillan 2002; Portegies Zwart et al. 2004, 2005; Giersz et al. 2015; Mapelli 2016; Spera & Mapelli 2017; Costa et al. 2023).
If VMSs do not undergo direct collapse at the end of their life but instead explode as supernovae, they release enormous amounts of material into the interstellar medium (Ohkubo et al. 2006; Yokoyama & Tominaga 2012; Kojima et al. 2021; Goswami et al. 2022; Vink 2023). Specifically, one of the possible final fates of VMSs is a pair-instability supernova (PISN) (Heger & Woosley 2002; Woosley 2017; Spera & Mapelli 2017). PISNe occur when the efficient formation of electron-positron pairs softens the equation of state of the carbon-oxygen core (Barkat et al. 1967; Fraley 1968; Bond et al. 1984; Heger & Woosley 2002). The resulting contraction of the core triggers explosive burning of oxygen and silicon. As a consequence, the star may undergo a single, powerful explosion that leaves no compact remnant (PISN), or it may experience multiple pulsations before ultimately collapsing into a black hole (pulsational pair-instability supernova, PPISN; Woosley et al. 2002; Chen et al. 2014; Yoshida et al. 2016; Woosley 2017). Since the onset of the instability depends on the central density and temperature, the final mass of the helium core ($M_{\mathrm{He,f}}$) is a good proxy to infer whether a star enters the pair-instability regime or not: according to Woosley (2017), if 32 $32\text{ M}_{\odot}\leq M_{\mathrm{He,f}}<64\text{ M}_{\odot}$, it undergoes a PPISN; if 64 $64\text{ M}_{\odot}\leq M_{\mathrm{He,f}}<135\text{ M}_{\odot}$, it explodes in a PISN; otherwise, if $M_{\mathrm{He,f}}>135\text{ M}_{\odot}$, it collapses directly onto an IMBH.
@@
Tanikawa et al. (2023, hereafter T23) explored the influence of the 12C($\alpha$,$\gamma$)16O reaction rate on the PISN rate in models of binary populations. By adopting the standard reaction rate and assuming a maximum zero-age main-sequence (ZAMS) mass of 300 M, they found $\mathcal{R}_{\mathrm{PISN}}\geq 1\mathrm{\,yr}^{-1}\mathrm{Gpc}^{-3}$. In contrast, using a $3\sigma$ lower reaction rate, they found $\mathcal{R}_{\mathrm{PISN}}\ll 1\mathrm{\,yr}^{-1}\mathrm{Gpc}^{-3}$. The critical role of the 12C($\alpha$,$\gamma$)16O reaction rate has been largely debated in the context of the pair-instability mass gap of black holes, with lower rates leading to higher black hole masses below the gap (e.g., Farmer et al. 2019, 2020; Costa et al. 2021; Woosley & Heger 2021).
@@
In summary, most theoretical estimates of the PISN rate density exceed the upper limit reported by Schulze et al. (2024). The only models that yield PISN rates below this upper limit assume low values for the 12C($\alpha$,$\gamma$)16O reaction rate, enforce a low metallicity threshold for the onset of PISNe, suppress the low-metallicity tails of star formation in the nearby Universe, and/or put a cap on the maximum ZAMS mass. The upper limit to the maximum ZAMS mass is particularly critical, because we have observed stars with $M_{\textsc{ZAMS}}>150$ M even in the local Universe (Crowther et al. 2010; Bestenlehner et al. 2014, 2020).
Here, we explore an alternative scenario inspired by recent models of optically thick winds for VMSs (Sabhahit et al. 2023, hereafter S23). This scenario does not require to fine tune the 12C($\alpha$,$\gamma$)16O reaction rate, works for a realistic evolution of the metallicity spread in galaxies, and does not require to cap the maximum star mass, while it naturally leads to a low metallicity threshold $Z_{\mathrm{th}}$ for the activation of PISNe ($Z_{\mathrm{th}}\sim{Z_{\odot}/10}$, S23).
Recent models of mass loss by stellar winds (Vink et al. 2011; Sabhahit et al. 2022, 2023) indicate that VMSs develop optically thick winds already on the main sequence, becoming WNh stars (Bestenlehner et al. 2020). The transition between optically thin winds of O-type stars and optically thick ones is a function of the electron-scattering Eddington ratio $\Gamma_{\mathrm{Edd}}$. At the metallicity of the Large Magellanic Cloud and even the Small Magellanic Cloud, this model strips the H-rich envelope of VMSs completely, without the need for mass loss through binary interaction. Here, we investigate the impact of such new stellar wind models (S23) on the expected PISN and PPISN rate density. We show that a larger mass-loss rate for VMSs at intermediate metallicity (down to $Z\sim{}0.002$, S23) seems to be the key to predicting PISN rates in agreement with the observational constraints.
@@ 2 Method
We use the stellar evolution code mesa (version r12115; Paxton et al. 2011, 2013, 2015, 2018, 2019) to model massive stars and VMSs and produce stellar tables for the population synthesis code sevn (Spera & Mapelli 2017; Spera et al. 2019; Mapelli et al. 2020; Iorio et al. 2023). Then, we combine our population synthesis models with a metal-dependent star formation rate density history by means of the semi-analytic code $\mathcal{R}$ate (Santoliquido et al. 2021, 2022). Here below, we describe our methodology in detail.
@@ 2.1 Stellar models with mesa
where $\dot{M}_{\mathrm{tr}}$, $L_{\mathrm{tr}}$ and $M_{\mathrm{tr}}$ are the mass-loss rate, the luminosity and the mass at the transition point. To obtain $\Gamma_{\mathrm{Edd,tr}}$, the following condition is iteratively verified:
@@
During the core He-burning phase, we use Eq. 1 only for temperatures between $4\times 10^{3}\mathrm{K}$ and $10^{4}$ K, while for lower temperatures ($T<4\times 10^{3}\mathrm{K}$), we follow de Jager et al. (1988). For $T>10^{4}$ K and hydrogen mass fraction $X<0.01$, we adopt the Wolf-Rayet (WR) mass-loss models by Sander & Vink (2020).
@@
We calculate stellar models with initial masses ranging from 50 to 500 M by intervals of 25 M. The initial metallicities are $Z$ = 0.008, 0.006, 0.004, 0.002, 0.001, 0.0008, 0.0006, 0.0004, 0.0002, and 0.0001. Thus, our grid consists of 190 models (19 initial masses and 10 initial metallicities). The initial helium mass fraction, $Y$, is computed as $Y=Y_{\mathrm{prim}}+(\Delta Y/\Delta Z),Z$, where $Y_{\mathrm{prim}}$ is the primordial helium abundance and $\Delta Y/\Delta Z$ describes the helium-to-metal ratio. We adopt the default values from mesa (Pols et al. 1998), with $Y_{\mathrm{prim}}=0.24$ and $\Delta Y/\Delta Z=2$. This formulation spans the range from a nearly primordial composition ($X=0.76$, $Y=0.24$, $Z=0$) to a quasi-solar one ($X=0.70$, $Y=0.28$, $Z=0.02$). The hydrogen mass fraction is given by $X=1-Y-Z$.
@@ 2.2 Population synthesis with sevn
We consider the same initial metallicities as in the mesa models. We simulate $10^{5}$ stars for each metallicity. The initial masses are distributed following a Kroupa initial mass function (IMF, Kroupa 2001) with slope $\alpha=2.3$ between 50 M and 500 M. To model pair-instability and pulsational pair-instability supernovae, we use the fitting formulas described by Mapelli et al. (2020), based on the models by Woosley (2017). Namely, a star undergoes a PPISN if the final mass of the He core is between 32 $32\text{ M}_{\odot}\leq M_{\mathrm{He,f}}<64\text{ M}_{\odot}$. A PISN takes place if 64 $64\text{ M}_{\odot}\leq M_{\mathrm{He,f}}\leq 135\text{ M}_{\odot}$. For higher masses of the He core, the star directly collapses onto an intermediate-mass BH. While a PISN leaves no compact remnant, the mass of the BH after a PPISN $M_{\mathrm{PPISN}}$ is estimated as
$M_{\mathrm{BH}}=\alpha_{\mathrm{P}}\,M_{\mathrm{CCSN}},$ (3)
where $M_{\mathrm{CCSN}}$ is the expected BH mass based on the fitting formulas by Fryer et al. (2012) and $\alpha_{\mathrm{P}}$ is a correction factor depending on the mass of He-core and on the ratio between the He-core mass and the total mass in the pre-supernova stage.
@@
We model the evolution of the (P)PISN rate density as a function of redshift with the code $\mathcal{R}$$\mathcal{R}$ate is publicly available at this gitlab repository. (Santoliquido et al. 2020, 2021), based on the observed metal-dependent cosmic star formation rate density (Madau & Fragos 2017).
The default version of $\mathcal{R}$ate estimates the merger rate density of compact binaries as a function of the redshift, but it can be adapted to trace the rate density of other transients of stellar origin, including (P)PISNe. Specifically, we calculate the rate of (P)PISNe as follows:
@@
where $\mathcal{S}(z^{\prime},Z)=\psi(z^{\prime})\,{}p(z^{\prime},Z)$.
Here, $\psi(z^{\prime})$ is the cosmic star formation rate density at redshift $z^{\prime}$, and $p(z^{\prime},Z)$ is the distribution of metallicity $Z$ at fixed formation redshift $z^{\prime}$. In Equation 4, ${{\rm{d}}t(z^{\prime})}/{{\rm{d}}z^{\prime}}=H_{0}^{-1}\,{}(1+z^{\prime})^{-1}\,{}[(1+z^{\prime})^{3}\Omega_{\mathrm{M}}+\Omega_{\Lambda}{}]^{-1/2}$, where $H_{0}$ is the Hubble parameter, $\Omega_{\mathrm{M}}$ and $\Omega_{\Lambda}$ are the matter and energy density, respectively. We adopt the values from Planck Collaboration et al. (2020).
@@
where ${{\rm{d}}\mathcal{N}(z^{\prime},z,Z)/{\rm{d}}}t(z)$ is the rate of (P)PISNe from stars with metallicity $Z$ that formed at redshift $z^{\prime}$ and undergo a (P)PISN at redshift $z$, and $M_{\ast}(z^{\prime},Z)$ is the initial total stellar mass that forms at redshift $z^{\prime}$ with metallicity $Z$. We obtain ${{\rm{d}}\mathcal{N}(z^{\prime},z,Z)/{\rm{d}}}t(z)$ directly from our simulations.
@@
The metallicity-evolution term $p(z^{\prime},Z)$ is the probability density function that a star formed at redshift $z^{\prime}$ has a metallicity $Z$, modelled as a log-normal distribution function (Chruslinska et al. 2019; Santoliquido et al. 2023):
$p(z^{\prime},Z)=\frac{1}{\sqrt{2\,\pi\,\sigma_{\mathrm{Z}}^{2}}}\,\exp{\left\{-\,\frac{\left[\log(Z/Z_{\odot})-\langle{}\log(Z/Z_{\odot})\rangle{}\right]^{2}}{2\,\sigma_{\mathrm{Z}}^{2}}\right\}}\,.$ (7)
@@
Figure 2: Contour plot showing the He core masses at the end of the He burning for our mesa tracks adopting S23 winds as a function of the metallicity and ZAMS mass. The white lines highlight the levels at $32\text{M}_{\odot}$, 64 $64\text{ M}_{\odot}$ and 135 $135\text{ M}_{\odot}$, which set the boundaries to have PPISNe, PISNe, and direct collapse via photodisintegration.

### v2

@@
Filippo Simonato e-mail: filippo.simonato@gssi.it112233    Stefano Torniamenti e-mail: sttorniamenti@mpia.de4455    Michela Mapelli e-mail: mapelli@uni-heidelberg.de1122556677 0000-0001-8799-2548    Giuliano Iorio 88 0000-0003-0293-503X    Lumen Boco 55    Franca De Domenico-Langer 55    Cecilia Sgalletta 9910101111 0009-0003-7951-4820
@@
Very massive stars (VMSs, $M_{\star}$ $\geq$ 100 M) play a crucial role in several astrophysical processes. At low metallicity, they might collapse directly into black holes, or end their lives as pair-instability supernovae. Recent observational results set an upper limit of $0.7\,{}\mathrm{yr}^{-1}\,{}\mathrm{Gpc}^{-3}$ on the rate density of pair-instability supernovae in the nearby Universe. However, most theoretical models predict rates exceeding this limit. Here, we compute new VMS tracks with the mesa code, and use them to analyze the evolution of the (pulsational) pair-instability supernova rate density across cosmic time. We show that stellar wind models accounting for the transition between optically thin and thick winds yield a pair-instability supernova rate $\mathcal{R}_{\mathrm{PISN}}\sim{}0.1$ Gpc-3 yr-1 at redshift $z\sim{}0$, about two orders of magnitude lower than our previous models. We find that the main contribution to the pair-instability supernova rate comes from stars with metallicity $Z\sim{}0.001-0.002$. Stars with higher metallicities cannot enter the pair-instability supernova regime, even if their zero-age main sequence mass is up to 500 M. The main reason is that VMSs enter the regime for optically thick winds during the main sequence at metallicity as low as $Z\sim{4}\times{}10^{-4}$. This enhances the mass loss rate, quenching the growth of the He core and thus preventing the onset of pair-instability in later evolutionary stages. This result highlights the critical role of mass loss in shaping the final fate of very massive stars and the rate of pair-instability supernovae.
@@ 1 Introduction
Very massive stars (VMSs), with an initial mass $100\text{ M}_{\odot}$, represent a challenge to our understanding of stellar evolution due to their extreme physical properties. Several studies have explored their evolutionary paths (Belkus et al. 2007; Yungelson et al. 2008; Gräfener et al. 2011; Brott et al. 2011; Georgy et al. 2013; Yusof et al. 2013; Hirschi 2015; Limongi & Chieffi 2018) and their observational properties in the local Universe (Crowther et al. 2010; Bestenlehner et al. 2014).
As VMSs evolve, they lose mass via stellar winds, including nitrogen and other products of hot hydrogen burning. Hence, they might be one of the main sources of the N enhancement observed in some compact high-redshift galaxies (Vink 2023; Cameron et al. 2023; Charbonnel et al. 2023; Gieles et al. 2025). Several works show that, if they directly collapse onto a black hole, VMSs can form intermediate-mass black holes (black holes with mass $10^{2}$ M, Ebisuzaki et al. 2001; Heger & Woosley 2002; Portegies Zwart & McMillan 2002; Portegies Zwart et al. 2004, 2005; Giersz et al. 2015; Mapelli 2016; Spera & Mapelli 2017; Costa et al. 2023).
If VMSs do not undergo direct collapse at the end of their life but instead explode as supernovae, they release enormous amounts of material into the interstellar medium (Ohkubo et al. 2006; Yokoyama & Tominaga 2012; Kojima et al. 2021; Goswami et al. 2022; Vink 2023). Specifically, one of the possible final fates of VMSs is a pair-instability supernova (PISN) (Heger & Woosley 2002; Woosley 2017; Spera & Mapelli 2017). PISNe occur when the efficient formation of electron-positron pairs softens the equation of state of the carbon-oxygen core (Barkat et al. 1967; Fraley 1968; Bond et al. 1984; Heger & Woosley 2002). The resulting contraction of the core triggers explosive burning of oxygen and silicon. As a consequence, the star may undergo a single, powerful explosion that leaves no compact remnant (PISN), or it may experience multiple pulsations before ultimately collapsing into a black hole (pulsational pair-instability supernova, PPISN; Woosley et al. 2002; Chen et al. 2014; Yoshida et al. 2016; Woosley 2017). Since the onset of the instability depends on the central density and temperature, the final mass of the helium core ($M_{\mathrm{He,f}}$) is a good proxy to infer whether a star enters the pair-instability regime or not: according to Woosley (2017), if $32\text{ M}_{\odot}\leq M_{\mathrm{He,f}}<64\text{ M}_{\odot}$, it undergoes a PPISN; if $64\text{ M}_{\odot}\leq M_{\mathrm{He,f}}<135\text{ M}_{\odot}$, it explodes in a PISN; otherwise, if $M_{\mathrm{He,f}}>135\text{ M}_{\odot}$, it collapses directly onto an intermediate-mass BH.
@@
Tanikawa et al. (2023, hereafter T23) explored the influence of the 12C($\alpha$$\gamma$)16O reaction rate on the PISN rate in models of binary populations. By adopting the standard reaction rate and assuming a maximum zero-age main-sequence (ZAMS) mass of 300 M, they found $\mathcal{R}_{\mathrm{PISN}}\geq 1\mathrm{\,yr}^{-1}\mathrm{Gpc}^{-3}$. In contrast, using a $3\sigma$ lower reaction rate, they found $\mathcal{R}_{\mathrm{PISN}}\ll 1\mathrm{\,yr}^{-1}\mathrm{Gpc}^{-3}$. The critical role of the 12C($\alpha$$\gamma$)16O reaction rate has been largely debated in the context of the pair-instability mass gap of black holes, with lower rates leading to higher black hole masses below the gap (e.g., Farmer et al. 2019, 2020; Costa et al. 2021; Woosley & Heger 2021).
@@
In summary, most theoretical estimates of the PISN rate density exceed the upper limit reported by Schulze et al. (2024). The only models that yield PISN rates below this upper limit assume low values for the 12C($\alpha$$\gamma$)16O reaction rate, enforce a low metallicity threshold for the onset of PISNe, suppress the low-metallicity tails of star formation in the nearby Universe, and/or put a cap on the maximum ZAMS mass. The upper limit to the maximum ZAMS mass is particularly critical, because we have observed stars with $M_{\textsc{ZAMS}}>150$ M even in the local Universe (Crowther et al. 2010; Bestenlehner et al. 2014, 2020).
Here, we explore an alternative scenario inspired by recent models of optically thick winds for VMSs (Sabhahit et al. 2023, hereafter S23). This scenario does not require to fine tune the 12C($\alpha$$\gamma$)16O reaction rate, works for a realistic evolution of the metallicity spread in galaxies, and does not require to cap the maximum star mass, while it naturally leads to a low metallicity threshold $Z_{\mathrm{th}}$ for the activation of PISNe ($Z_{\mathrm{th}}\sim{Z_{\odot}/10}$, S23).
Although the stellar winds of massive and very massive stars remain an active topic of research (Yang et al. 2023; Cheng et al. 2024; Pauli et al. 2025), recent models of mass loss by stellar winds (Vink et al. 2011; Sabhahit et al. 2022, 2023) indicate that VMSs develop optically thick winds already on the main sequence, becoming WNh stars (Bestenlehner et al. 2020). The transition between optically thin winds of O-type stars and optically thick ones is a function of the electron-scattering Eddington ratio $\Gamma_{\mathrm{Edd}}$. At the metallicity of the Large Magellanic Cloud and even the Small Magellanic Cloud, this model strips the H-rich envelope of VMSs completely, without the need for mass loss through binary interaction. Here, we investigate the impact of such new stellar wind models (S23) on the expected PISN and PPISN rate density. We show that a larger mass-loss rate for VMSs at intermediate metallicity (down to $Z\sim{}0.002$, S23) seems to be the key to predicting PISN rates in agreement with the observational constraints.
@@ 2 Method
[…]
@@ 2.1 Stellar models with mesa
where $\dot{M}_{\mathrm{tr}}$, $L_{\mathrm{tr}}$, and $M_{\mathrm{tr}}$ are the mass-loss rate, the luminosity, and the mass at the transition point. To obtain $\Gamma_{\mathrm{Edd,tr}}$, the following condition is iteratively verified:
@@
During the core He-burning phase, we use Eq. 1 only for temperatures between $4\times 10^{3}\mathrm{K}$ and $10^{5}$ K, while for lower temperatures ($T<4\times 10^{3}\mathrm{K}$), we follow de Jager et al. (1988). For $T>10^{5}$ K and hydrogen mass fraction $X_{\text{C}}<0.01$, we adopt the Wolf-Rayet (WR) mass-loss models by Sander & Vink (2020). Recently, Boco et al. (2025) showed that this formalism, when combined with fast rotation, can strip single metal-poor O-type stars and convert them into WNh stars even at low metallicity ($Z\sim{0.0002}$). Thus, this mechanism offers a potential explanation for the presence of WR stars in metal-poor galaxies, such as the Small Magellanic Cloud (Schootemeijer et al. 2021).
@@
We calculate stellar models with initial masses ranging from 50 to 500 M by intervals of 25 M. The initial metallicities are $Z$ = 0.008, 0.006, 0.004, 0.002, 0.001, 0.0008, 0.0006, 0.0004, 0.0002, and 0.0001. Thus, our grid consists of 190 models (19 initial masses and 10 initial metallicities). The initial helium mass fraction, $Y$, is computed as $Y=Y_{\mathrm{prim}}+(\Delta Y/\Delta Z),Z$, where $Y_{\mathrm{prim}}$ is the primordial helium abundance and $\Delta Y/\Delta Z$ describes the helium-to-metal ratio. We adopt the default values from mesa (Pols et al. 1998), with $Y_{\mathrm{prim}}=0.24$ and $\Delta Y/\Delta Z=2$. This formulation spans the range from a nearly primordial composition ($X=0.76$, $Y=0.24$, $Z=0$) to a quasi-solar one ($X=0.70$, $Y=0.28$, $Z=0.02$). The hydrogen mass fraction is given by X=1YZX=1-Y-Z.
@@ 2.2 Population synthesis with sevn
We consider the same initial metallicities as in the mesa models. We simulate $10^{5}$ stars for each metallicity. The initial masses are distributed following a Kroupa initial mass function (IMF, Kroupa 2001) with slope $\alpha=2.3$ between 50 M and 500 M. To model pair-instability and pulsational pair-instability supernovae, we use the fitting formulas described by Mapelli et al. (2020), based on the models by Woosley (2017). Namely, a star undergoes a PPISN if the final mass of the He core is between $32\text{ M}_{\odot}\leq M_{\mathrm{He,f}}<64\text{ M}_{\odot}$. A PISN takes place if $64\text{ M}_{\odot}\leq M_{\mathrm{He,f}}\leq 135\text{ M}_{\odot}$. For higher masses of the He core, the star directly collapses onto an intermediate-mass BH.
@@
[…]
@@
where $\mathcal{S}(z^{\prime},Z)=\psi(z^{\prime})\,{}{p(Z|z^{\prime})}$.
Here, $\psi(z^{\prime})$ is the cosmic star formation rate density at redshift $z^{\prime}$, and $p(Z|z^{\prime})$ is the distribution of metallicity $Z$ at fixed formation redshift $z^{\prime}$. In Equation 3, ${{\rm{d}}t(z^{\prime})}/{{\rm{d}}z^{\prime}}=H_{0}^{-1}\,{}(1+z^{\prime})^{-1}\,{}[(1+z^{\prime})^{3}\Omega_{\mathrm{M}}+\Omega_{\Lambda}{}]^{-1/2}$, where $H_{0}$ is the Hubble parameter, $\Omega_{\mathrm{M}}$ and $\Omega_{\Lambda}$ are the matter and energy density, respectively. We adopt the values from Planck Collaboration et al. (2020).
@@
where ${{\rm{d}}\mathcal{N}(z^{\prime},z,Z)/{\rm{d}}}t(z)$ is the rate of (P)PISNe from stars with metallicity $Z$ that formed at redshift $z^{\prime}$ and undergo a (P)PISN at redshift $z$, and $M_{\ast}(z^{\prime},Z)$ is the initial total stellar mass that forms at redshift $z^{\prime}$ with metallicity $Z$. We obtain ${{\rm{d}}\mathcal{N}(z^{\prime},z,Z)/{\rm{d}}}t(z)$ directly from our simulations. Note that $\mathcal{N}(z^{\prime},z,Z)$ also depends on the stellar lifetime (hence on the stellar mass) and on the assumed initial mass function. While we do not explicitly write down these dependences for notation simplicity, our catalogs fully account for them.
@@
The metallicity-evolution term $p(Z|z^{\prime})$ is the probability density function that a star formed at redshift $z^{\prime}$ has a metallicity $Z$, modelled as a log-normal distribution function (Chruslinska et al. 2019; Santoliquido et al. 2023):
${p(Z|z^{\prime})}=\frac{1}{\sqrt{2\,\pi\,\sigma_{\mathrm{Z}}^{2}}}\,\exp{\left\{-\,\frac{\left[\log(Z/Z_{\odot})-\langle{}\log(Z/Z_{\odot})\rangle{}\right]^{2}}{2\,\sigma_{\mathrm{Z}}^{2}}\right\}}\,.$ (6)
@@
Figure 2: Contour plot showing the He core masses at the end of the He burning for our mesa tracks adopting S23 winds as a function of the metallicity and ZAMS mass. The white lines highlight the levels at $32\text{M}_{\odot}$, $64\text{ M}_{\odot}$ and $135\text{ M}_{\odot}$, which set the boundaries to have PPISNe, PISNe, and direct collapse via photodisintegration.